### v2

# Does the suppression shown at LHC in O-O collisions follow the systematics obtained for A-A collisions ??

Abstract
In this paper we present to what extent recent experimental results obtained for $\pi^{0}$ suppression in O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV fit into the systematics for much heavier systems. The systematics with which the comparison is made was published a few years ago [14] in terms of charged particles suppression $R_{AA}$ as a function of $\langle N_{part}\rangle$ and $\langle dN_{ch}/d\eta\rangle$. The values of $R_{AA}$ at $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ experimentally measured and estimated by Glauber MC, respectively, for O-O collisions are in good agreement with the systematics obtained for A-A collisions.
I Introduction
The studies of $p$ collisions at LHC energies up to very high charged particle multiplicities, revealed similarities between $p$ and Pb-Pb in the behaviour of different observables, like near-side long range pseudorapidity correlations [8], the ($\langle\beta_{T}\rangle$ - $T_{kin}^{fo}$) correlation as a function of charged particle multiplicity [6], azimuthal angular correlations [9], geometrical scaling [13, 15, 16], etc. Based on the suppression studies in Cu-Cu and Au-Au collisions at RHIC and Xe-Xe and Pb-Pb collisions at LHC and the dependence of $R_{AA}^{N}=\frac{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{cen}}{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp,INEL}}$ on $\langle dN_{ch}/d\eta\rangle$ [14], it was argued why even in high charged particle multiplicity events in $p$ collisions at LHC, in the limit of current experimental uncertainties, no suppression was observed, although similarities to Pb-Pb collisions for other observables have been highlighted. Recently, at the LHC energy of $\sqrt{s_{NN}}$=5.36 TeV, O-O collsions were measured and preliminary $R_{AA}$ results were reported by ALICE [18, 12] and CMS [7] Collaborations for $\pi^{0}$ and charged particles, respectively. In this letter we are comparing the ALICE results with the results of a systematic study of $R_{AA}$ as a function of $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ for A-A collisions.
II $\langle N_{part}\rangle$ dependence of $\langle N_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{O-O}/\langle dN_{ch}/d\eta\rangle^{pp}$]
In the paper cited above [14] was also presented a dependence on $\langle N_{part}\rangle$ of $\langle N_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] for Cu-Cu and Au-Au at $\sqrt{s_{NN}}$=200 GeV, Xe-Xe at $\sqrt{s_{NN}}$=5.44 TeV, Pb-Pb at $\sqrt{s_{NN}}$=2.76 TeV and 5.02 TeV using for $\langle dN_{ch}/d\eta\rangle$ the values reported in [2, 5, 3, 1, 4], $N_{bin}>$ estimated within the Glauber MC approach [10] and the $\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp}$ INEL values at the corresponding collision energies (Fig.1). Within the error bars, a good scaling is observed. The rather large values of the error bars are due to uncertainty in $\langle N_{bin}\rangle$ estimates [10]. In Fig.1 are represented by yellow squares the values corresponding to O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV for different centralities, using $\langle dN_{ch}/d\eta\rangle$ reported by the ALICE Collaboration [18, 12]. $\langle N_{part}\rangle$ and $\langle N_{bin}\rangle$ estimated within the Glauber MC approach were taken from [11] for the corresponding centralities. As can be observed, the values for O-O collisions follow the systematics obtained in heavy ion collisions.
Figure 1: $\langle dN_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] values in O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV are presented by yellow squares using recently measured $\langle dN_{ch}/d\eta\rangle$ for different centralities by the ALICE Collaboration [18, 12] on top of Fig.14 from [14]
Figure 2: $R_{AA}$ as a function of $\langle N_{part}\rangle$ from [13] over which the corresponding experimental value for O-O is represented by the yellow box.
III $R_{AA}$ - $\langle N_{part}\rangle$ dependence
As it was presented in a previous paper [13], the transverse overlap area of the nucleons suffering more that single collisions $S_{\perp}^{core}$ shows a very good scaling as a function of $N_{part}$, both being estimated within the Glauber MC approach for the Au-Au at energies measured at RHIC and Pb-Pb at energies measured at LHC. Similar results were obtained for $S_{\perp}^{var,core}$ for Cu-Cu, Au-Au at top energy at RHIC and Xe-Xe and Pb-Pb at energies measured at LHC. The aspect ratios of the transverse overlap area $S_{\perp}$ for different systems and energies are also rather similar. With the assumption that the average path length of a parton traveling the deconfined zone $L^{2}\propto S_{\perp}$ one could expect a good scaling of the average suppression as a function of $N_{part}$ at a given collision energy. This was presented in ref. [14]. For O-O at $\sqrt{s_{NN}}$=5.36 TeV, an estimate of $N_{part}$=10.8 for 0-100% centrality based on the Glauber MC approach was reported in a recent publication [11]. In Fig.2 by the yellow box is represented the $R_{AA}$ value reported as preliminary results by the ALICE Collaboration [18, 12] at the corresponding $N_{part}$ on top of the compilation presented in [14]. The agreement between the O-O experimental result and the $N_{part}$ dependence of $R_{AA}$ obtained in Xe-Xe and Pb-Pb at similar collision energies is very good.
IV $R_{AA}-\langle dN_{ch}/d\eta\rangle$ dependence
Within the phenomenological parton energy loss approach, the fractional energy loss is given by:
$\frac{\Delta E}{E}\propto T^{a}L^{b}$ (1)
where a and b take different values [20]. This could explain the collision energy dependence of $R_{AA}$ as a function of $N_{part}$, the temperature being rather different as far as $\langle dN_{ch}/d\eta\rangle$ is different for a given value of $N_{part}$ [14]. Based on this argument, in [14] we also presented $R_{AA}$ as a function of $dN_{ch}/d\eta$ for the same systems. In such a representation the energy dependence of the scaling does not appear anymore although the aspect ratio of $S_{\perp}$ for different systems and collision energy is rather different. The study of the azimuthal dependence of $R_{AA}$ done by the PHENIX Collaboration for Au-Au collision at $\sqrt{s_{NN}}$=200 GeV and 30-40% centrality showed a larger suppression out of the reaction plane relative to the one in-plane [17]. It seems that this effect is less important in the average value of $R_{AA}$ and the temperature dependence of the energy loss in the deconfined matter plays the main role.
Figure 3: $R_{AA}$ as a function of $\langle dN_{ch}/d\eta\rangle$ from [13] on which the corresponding experimental value for O-O is represented by the yellow box.
Taking the measured value of $\langle dN_{ch}/d\eta\rangle$ for O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV reported by the ALICE Collaboration [18, 12] , the result is presented in Fig.3 by the yellow square on top of the compilation presented in Fig.8 of reference [13]. The $R_{AA}$ for the O-O case is in good agreement with the results obtained for much heavier systems at RHIC and LHC at the same $\langle dN_{ch}/d\eta\rangle$.
V Conclusions
Although the alpha-like cluster structure of the colliding O nuclei could result in peculiar behaviour of many observables, it seems that the average value of $R_{AA}$ is in very good agreement with those extracted from Pb-Pb and Xe-Xe collisions at a similar collision energy at the same $\langle dN_{ch}/d\eta\rangle$ or $\langle N_{part}\rangle$ values. The expected results from detailed multi differetial data analysis of the information collected in alpha-like light ion collisions at LHC will show to which extent the previously established systematics in heavy ion collisions and $p$ at LHC energies [14, 15, 16, 19] are also followed in the case of such collisions.